### v1

# IQA-Eval: Automatic Evaluation of Human-Model Interactive Question Answering

Abstract
To evaluate Large Language Models (LLMs) for question answering (QA), traditional methods typically focus on directly assessing the immediate responses generated by the models based on the given question and context. In the common use case of humans seeking AI assistant’s help in finding information, these non-interactive evaluations do not account for the dynamic nature of human-model conversations, and interaction-aware evaluations have shown that accurate QA models are preferred by humans 24. Recent works in human-computer interaction (HCI) have employed human evaluators to conduct interactions and evaluations, but they are often prohibitively expensive and time-consuming to scale. In this work, we introduce an automatic evaluation framework IQA-Eval to Interactive Question Answering Evaluations [^1] https://github.com/du-nlp-lab/IQA-Eval, more specifically, we introduce LLM-based Evaluation Agent (LEA) that can: (1) simulate human behaviors to generate interactions with IQA models; (2) automatically evaluate the generated interactions. Moreover, we propose assigning personas to LEAs to better simulate groups of real human evaluators. We show that: (1) our evaluation framework with GPT-4 (or Claude) as the backbone model achieves a high correlation with human evaluations on the IQA task; (2) assigning personas to LEA to better represent the crowd further significantly improves correlations. Finally, we use our automatic metric to evaluate five recent representative LLMs with over 1000 questions from complex and ambiguous question answering tasks, which comes with a substantial cost of \$5k if evaluated by humans.
1 Introduction
The advent of Large Language Models (LLMs) has significantly advanced the field of natural language processing (NLP), enabling systems to perform a wide range of tasks with remarkable proficiency 55, 48, 5. Among these tasks, question answering (QA) has emerged as a critical and representative goal-oriented application, demonstrating the potential of LLMs to generate informative responses as an assistant 4. The success of these applications or in general the success of LLMs, heavily relies on accurate and consistent evaluation 8. Accurate and consistent evaluation helps researchers better understand LLMs capacities, and offers insights for new human-LLMs QA interactions 29, 8.
Figure 1: An example of human-model interactive question answering (IQA) and our automatic evaluation (IQA-Eval). The two interactions occur with two types of personas in humans (or LLM-based evaluation agents): Expert and Clarity-seeker, and are evaluated by humans or agents with corresponding personas. The IQA model only responds to the immediately preceding prompt without further contexts like the question itself (leftmost in the Figure).
Traditionally, automatic metrics such as accuracy have been used to evaluate models based on the quality of their direct answers to specific questions. However, as interactions between humans and LLMs grow more complex and nuanced, these traditional metrics often fail to capture the full spectrum of a model’s capabilities (e.g. helpfulness and providing enjoyment), particularly in interactive QA settings 30, 12, whereas interactions are crucial for user experience and system effectiveness, yet remains overlooked in traditional evaluation paradigms. Recently work such as 24 evaluated human-LLM interactions, a process that involves human participation and annotation. Although human evaluations for these interactions provide a closer approximation to real-world use cases, this approach is significantly costly and time-consuming and could also be prone to errors. Recognizing the need for automatic evaluation, works like G-Eval 32, LLM-Eval 29 and PRD 27 proposed to automate the assessment of non-interactive LLM responses using LLMs.
Drawing insights from (a) using LLM for automatic evaluation and (b) literature of LLM-agents research 46, 13, we propose IQA-Eval framework to auto-evaluate the performance of IQA models (i.e. LLMs) with LLM-based Evaluation Agent (LEA) to simulate and then evaluate interactions. By incorporating personas, our experiments on a well-annotated dataset show that our methods align well with human judgments and provide a more comprehensive evaluation of LLMs in interactive settings than traditional metrics. See an overview of our system in Figure 1.
Meta-evaluation shows that IQA-Eval results highly correlate with human evaluations across metrics and in general (with scores being over 0.6). The well-alignment with human judgments underscores IQA-Eval as a significant advancement in the field of automated IQA evaluation, offering a robust alternative to traditional/manual methods. Finally, we benchmark recent LLMs with new complex and ambiguous questions, and find that the accuracy of answers does not transfer to the corresponding ranking of models on their capability of achieving good human-model interactions.
Our contributions are as follows:
  • We propose the first LLM agent-based automatic evaluation framework IQA-Eval  designed specifically to generate and then evaluate interactions in IQA. Our results demonstrate a strong correlation with human evaluations.
  • We propose assigning personas to the LLM evaluation agent, enabling a more nuanced and accurate assessment of a model’s ability to adapt its responses to real-world interaction scenarios, where the context and individual user preference significantly influence the nature of the interactions.
  • We experiment with IQA-Eval framework to benchmark the most recent LLMs on the IQA task, demonstrating the strength and general effectiveness of our framework on fully automated evaluation.
2 Related Work
Evaluating Interactions
Traditional methods for human-model dialogue evaluation have often been centered around single-turn pairwise evaluation 44, 25. Some methods with multi-turn Likert scores emerge 43, 54, 40 but require time-consuming and costly collection of human-model conversations. In response, 16 suggests a self-play scenario where a dialog system engages in conversation with itself, employing a set of proxy measures for evaluation.
Acute-eval 26 and LLM-Eval 29 have advanced multi-turn evaluation frameworks, reflecting the increasing demand for sophisticated techniques that holistically capture human-AI interactions. Further studies in specific interaction domains like task completion 31, code generation 52, and collaborative problem-solving 24, 19, 14 emphasize the need for evaluations that consider both environmental and human elements 47. Our work differs from these methods by introducing an automated approach that emphasizes interaction quality and significantly reduces the reliance on human annotations.
Personas in NLP
Personas are constructed profiles that represent key traits of a group of users, as defined in the HCI field, reflecting their characteristics, behaviors, and goals to guide the design of technologies that are well-suited to user needs 11. This approach is a promising direction for enhancing relevance and personalization in NLP applications 36, 2, 41, 57, offering significant potential for customizing engagement and improving the effectiveness of conversational agents 25, 54, wolf2019transfertransfo, 7, 33, 56. Li et al. 25 introduces a persona-based neural conversation model to enhance dialogue personalization and coherence. Zhang et al. 54 develops personalized dialogue agents that incorporate user-specific details to enhance interaction. In our work, persona settings enable our framework to tailor interactions and assessments, aligning more closely with the specific characteristics and preferences of different user groups.
LLM-based Agent for Simulation
Recently, LLMs demonstrated human-like intelligence, as evidenced in various studies 51, 39, 28, 21, 3, 6, 42, 10. The integration of LLMs into agents that simulate complex human behaviors and social interactions is an area of growing research interest 34, 49, 50. For example, Park et al. 37 and Gao et al. 15 employ these agents to simulate human emotions, attitudes, and behaviors in social networks, while Park et al. 38 leverages in-context learning to simulate human behaviors in sandbox world. Moreover, Argyle et al. 1 utilizes ”algorithmic bias” inherent in GPT-3 to reflect the response patterns of different human subgroups. Horton 18 utilize LLMs in experimental setups of behavioral economics experiments to facilitate pilot studies. Additionally, Hämäläinen et al. 20 and Wang et al. 45 investigate LLM-based agents in recommender systems to simulate and collect data on user behavior. These studies show the broad applicability and potential of LLM-based agents in simulating human behaviors and interactions across diverse applications. Our work utilizes LLM-based evaluation agents(LEAs) to fully automate interactive quality assessments, handling both interaction generation and evaluation, to enhance the evaluation of IQA models in realistic scenarios.
3 IQA-Eval: Evaluating Interactive Question Answering (IQA)
In this section, we introduce our IQA-Eval framework for automatically evaluating Interaction Question Answering Models (IQA models) with LLM-based Evaluation Agents (LEA). IQA-Eval utilizes LEA to simulate humans in the following two stages: (1) generating interactions with IQA models; and (2) evaluating interactions. Additionally, we explore assigning personas to the IQA-Eval to better simulate the diversity of the groups of humans and provide individualized evaluations.
3.1 Interaction Generation with LEA (Stage 1)
Inspired by peer discussions 24, 46, we propose utilizing LEAs with prompt-based techniques to simulate human behaviors for effective interaction generation with IQA models. The structured prompt includes three key components: (1) a role description; (2) a task description; and (3) instructions for the discussion.
Role Description outlines the people that the LEA model will simulate during interactions. For example, the description for a standard persona could be: You are mimicking a human.
Task Description briefly describes the action that LEA model needs to perform in the task. For example, in a multi-choice question answering task, the prompt could be structured as follows: You are trying to choose the correct answer for the given question. Both role and task descriptions can be adjusted based on the persona, as discussed in Section 3.3.
Discussion instruction guides LEAs on their subsequent steps by providing detailed descriptions to facilitate progress in interactions. It comprises two essential components: (1) the actions to take; and (2) the detailed procedures to follow. For example, in a question answering task, the prompt specifies:
You can ask an assistant questions for help. Please ask sub-questions
to approach answers. In each turn, please only ask one sub-question
to interact with the assistant.
In the sub-question, please include all necessary information in the
original question, such as the question and all options. If you know
the answer, please output "So, the answer is: A, B, C, or D".
The first paragraph of the discussion instruction outlines “the actions to take” and the second paragraph details “procedure to follow”.
At the start of an interaction, the LEA receives a system prompt that includes all three components above, along with the specific question to be addressed. As the LEA interacts with the IQA model, it generates sub-questions to request clarification of unknown entities, definitions, or particular aspects of the original question. Then, the IQA Model takes the questions as the input and output responses. After receiving responses, the LEA continues to pose further questions until it determines the final answer. The full prompt structure and interaction details are provided in Appendix B.2.
3.2 Interaction Evaluation with LEA (Stage 2)
Inspired by G-eval 32, which demonstrates that evaluations by GPT models align closely with human assessments in NLG tasks, we propose utilizing LEAs for interaction evaluation. This LEA assesses interactions generated by LEAs and IQAs in Stage 1. The module takes task details, such as questions or articles, and interactions as the input, and output evaluation scores. The prompt contains three parts: (1) role and task description; (2) metrics definition; and (3) evaluation instruction.
Role and task description instructs LEA to conduct evaluation. The role description acts the same as the role description in the previous stage. Moreover, it briefly describes the evaluation task. The general prompt looks like: You are a helpful and precise evaluator who checks the quality of the AI assistant’s responses in interactions.
Metrics definitions describes the criterion that LEA needs to follow in the evaluation process. They can customized for different tasks. For a question answering task, we add the following prompt to define the “helpfulness” metric:
Helpfulness (5-point Likert): How helpful was having access to the
AI Assistant compared to not having access?
Both the aforementioned parts can be tailored according to the persona that the LEA simulates in Stage 1, as detailed in Section 3.3.
Evaluation instruction outlines the specifics of the evaluation task and the required output format. This section may appear as a separate part of the prompt. For example, the instruction within the prompt might be structured as follows:
Please evaluate the above interactions between user and AI assistant
by using the following metrics:
<Metric definitions>
Please output each of the above metrics line-by-line.
Finally, all evaluation scores for metrics are calculated by averaging the results of multiple runs. The complete prompt is available in Appendix B.1, and further details about our implementation can be found in Section 4.
3.3 Assigning the Personas to LEA
The evaluation process for LLM IQA models, including interaction generation and interaction evaluation, typically utilizes a standard persona, designed to reflect neutral and average characteristics across various attributes such as knowledge, language proficiency, and beliefs. This approach ensures a consistent and uniform baseline for assessment. However, individual users often exhibit unique personal preferences and characteristics, making a one-size-fits-all evaluation less effective. Moreover, the persona distribution of the target user group significantly impacts the performance of IQA models in real-world applications. For instance, if 20% of human users prefer brief interactions and 70% prefer detailed information, applying a general LEA to simulate this group of persons is likely to result in poor correlation in terms of evaluations.
To better simulate the diversity of the groups of people and provide individualized evaluations, we assign personas to LEAs. This affects prompts of both interaction generation and interaction evaluation processes.
For example, when the LEA is assigned with the “Critical-Seeker” persona (definition in B.3) for interaction generation, we adapt the default role and task description (in Stage 1) to:
You prefer interactions rich in critical information. You need help from an assistant and try to get critical information from it to answer the following questions. For interaction evaluation, the default role and task description prompt changes to The AI Assistant should provide straightforward, simple, and concise answers to aid users in deducing solutions. Additionally, the definition of metrics is also adjusted to align with this persona, with further details available in Appendix B.3.
4 Meta-Evaluation of IQA-Eval Framework
To measure how our framework provides trustworthy IQA evaluations that align with human preferences, we conduct meta-evaluations experiments and report correlation scores.
4.1 Experiment Settings
Dataset and Evaluation Metrics
We apply our evaluation method on the annotated dataset from the study by Lee et al. 24. This dataset consists of 3641 interactions from 331 annotators. Questions in the dataset are multi-choice and are derived from the MMLU dataset 17 (example question in Figure 1). The construction of MMLU requires each worker to participate in 11 random conversations with one of the following three IQA models: TextDavinci (text-davinci-001), TextBabbage (text-babbage-001), and Davinci (davinci-001). At the end of conversations, fluency and helpfulness scores are annotated by annotators. The number of queries and accuracy for each IQA model can be easily deduced from annotations. In this work, We adjust the four metrics to evaluate generated interactions:
  • Fluency (5-point Likert): How clear (or fluent) were the responses from the AI Assistant?
  • Helpfulness (5-point Likert): Independent of its fluency, how helpful was having access to the AI Assistant compared to not having access?
  • Number of Queries: Counts the number of interaction turns in the conversation. This metric helps assess the efficiency of the AI in resolving queries within a minimal number of interactions.
  • Accuracy: Quantifies how accurately the AI’s responses match the golden answers. This is critical for evaluating the correctness of the AI’s knowledge and its application in practical scenarios.
LEA Models
To evaluate the effectiveness of IQA-Eval framework, we experiment with different LEA models on the above-mentioned three LLM IQA models in MMLU. For LEA that conducts both interaction generation and interaction evaluation, we use ChatGPT (GPT-3.5-turbo-1106), GPT4 (GPT-4-1106-preview), Claude (Claude-1).
Evaluation of IQA-Eval Framework
To evaluate the performance of our IQA-Eval framework, we use Pearson correlations as the indicator between the Human evaluations and LEA evaluations of IQA models.
4.2 Experiment Results
According to Table 2, all models, including GPT4, GPT3.5, and Claude, show high correlation with human evaluations. GPT4 aligns most closely with human judgments in both “Helpfulness” and “Fluency” metrics and the highest overall correlation score. This indicates that these models are capable of effectively performing IQA-Eval framework as LEA models.
Helpfulness Fluency # Queries Accuracy
Evaluator TDA TB DA TDA TB DA TDA TB DA TDA TB DA
Human 4.60 3.84 3.52 4.35 3.84 3.22 1.78 2.57 2.66 69.00 52.00 48.00
IQA-Eval-GPT4 3.67 2.30 2.10 4.77 3.87 3.03 1.57 2.27 2.37 0.87 0.83 0.67
IQA-Eval-Claude 4.13 3.03 3.00 4.47 3.47 3.23 2.20 2.67 2.07 0.67 0.53 0.57
IQA-Eval-GPT3.5 4.30 3.87 3.93 4.47 3.67 3.97 1.57 1.77 2.00 0.63 0.47 0.53
Table 1: IQA-Eval evaluation results of IQA models (TDA: TextDavinci; TB: TextBabbage; DA: Davinci). Bold numbers indicate they are the most close to human results. The empty set symbol (Ø) indicates the number cannot be calculated due to the model’s inability to follow instructions and produce a gradable answer.
Helpfulness Fluency Overall
IQA-Eval-GPT4 0.652 0.591 0.613
IQA-Eval-Claude 0.640 0.552 0.551
IQA-Eval-GPT3.5 0.621 0.523 0.510
Table 2: Pearson Correlation ($\rho$) between IQA-Eval evaluations and human judgments.
In Table 1, ChatGPT scores closest to human judgments, particularly in the ”Helpfulness” metric. Conversely, GPT4 and Claude score lower on ”Helpfulness” than human evaluations, because they tend to produce inaccurate and repetitive responses that lack coherence and do not directly address user queries, as indicated by their generated explanations. IQA-Eval scores on “Fluency” are close and highly correlated to human judgments. Both scores given by human and LEA models show that IQA models provide fluent outputs. Furthermore, according to the ”# Queries” metric, most models conclude conversations more quickly than humans, except for Claude, which requires more turns, potentially due to its non-OpenAI origins that it needs more turns to adapt the conversational style and understand responses. Notably, GPT4 achieves the highest accuracy among all models.
Analysis of LEA for Stage 2 (Evaluating Interactions)
For Stage 2 itself, we measure the LEA’s capability of evaluating interactions, based on real human-generated interactions from Stage 1.
(a) Helpfulness
(b) Fluency
Figure 2: Interaction evaluation results evaluated by human and two LEA models on interactions between real human and IQA Models. All scores are on a scale of 5.
Results are in shown Figures 2(a) and 2(b). The Pearson correlation coefficients for fluency and helpfulness between human judgments and LEA evaluations show distinct patterns. ChatGPT demonstrates a stronger correlation with human ratings, recording a correlation score of 0.424 for fluency and a correlation score of 0.306 for helpfulness. In contrast, Claude shows slightly lower correlations (0.281 for fluency and 0.287 for helpfulness. These results indicate that ChatGPT aligns better with human evaluative standards compared to Claude in these specific metrics.
Moreover, for interactions between humans and IQA models (Figure 2), LEA evaluations moderately correlate with human evaluations. However, for interactions between LEAs and IQA models (Table 2), which is the main focus of our paper, LEA evaluations highly correlate (around 0.6) with human evaluations. This indicates that LEA models are helpful when participating in the whole evaluation process, including both the interaction and evaluation. However, when evaluating interactions between human and IQA models, LEAs focus differently from humans, which causes moderate correlations.
Thus, the results show that (1) both models’ evaluations moderately correlate with human evaluation; (2) ChatGPT’s evaluation is closer and related to human evaluation than Claude’s.
4.3 Further Analysis for Free-form Feedback
Apart from the metrics above, we also prompt LLM-based Evaluation Agent (LEA) to explain the reason for generating scores in the free-form text format. We find that: 1) ChatGPT Generates more human-like, positive reviews: ChatGPT evaluations are generally more positive, frequently using terms like “helpful”, “relevant”, and “useful” – words not always noted by workers in their annotations. Despite this, ChatGPT often identifies similar issues as human raters, such as the provision of irrelevant and repetitive information. Overall, ChatGPT’s assessments align well with human evaluations; 2) Claude flags more Issues: Claude is more strict and critical to IQA models in interactions. In the free-text feedback, Claude tends to highlight more issues with model responses rather than acknowledging positive aspects, especially for Davinci. For one question, after interacting with the IQA model/assistant (ChatGPT in this case), human and two LEAs provide the following feedback:
5 Effect of Assigning Persona to LLM Evaluation Agent (LEA)
5.1 Persona Definitions
As discussed in Section 3.3, we assign personas to LEAs to simulate different groups of humans for diverse human alignments. We investigate assigning the following personas to LEAs, which are defined based on the crowdworker survey results in Lee et al. 24.
  • Expert: knowledgeable; quickly learns new concepts and applies them in the reasoning process to answer questions.
  • Critical-Thinker: people who prefer critical information rather than redundant or detailed responses.
  • Adaptability-Seeker: people who prefer assistants can understand their questions even if they are not precise.
  • Clarity-Seeker: people who prefer clear explanations from assistants.
Based on the survey results on the distributions of personas of workers 24, for each persona $P$, we split the crowdworkers into two groups: “persons with persona $P$”, and “normal persons without specific persona $P$”. For the first group, we initialize the role prompt in Section 3.3 and use it for the LEAs. For the second group, we utilize default prompting for LEAs. The LEA model in this section is ChatGPT(GPT-3.5-turbo-1106).
5.2 Experimental Results
# Queries Accuracy
Evaluator TDA TB TD TDA TB TD
Human 1.78 2.57 2.66 0.69 0.52 0.48
IQA-Eval 1.57 1.77 2.00 0.63 0.47 0.53
IQA-Eval (Expert) 1.20 1.49 2.20 0.73 0.56 0.53
IQA-Eval (Critical-Thinker) 1.55 1.80 1.99 0.68 0.54 0.55
IQA-Eval (Adaptability-Seeker) 1.50 1.75 2.10 0.66 0.52 0.55
IQA-Eval (Clarity-Seeker) 1.64 2.10 2.34 0.63 0.57 0.57
Table 3: IQA-Eval evaluation results of IQA models (TDA: TextDavinci; TB: TextBabbage; DA: Davinci). LEAs are assigned specific personas when representing specific groups of workers.
Helpfulness Fluency Overall
Evaluator TDA TB TD $\rho$ TDA TB TD $\rho$ $\rho$
Human 4.60 3.84 3.52 - 4.35 3.84 3.22 - -
IQA-Eval 4.30 (±0.06) 3.87 (±0.11) 3.93 (±0.13) 0.621 4.47 (±0.05) 3.67 (±0.08) 3.97 (±0.06) 0.523 0.510
IQA-Eval (Expert) 4.17 (±0.08) 3.08 (±0.09) 3.12 (±0.11) 0.756 4.47 (±0.02) 3.84 (±0.04) 3.40 (±0.04) 0.787 0.670
IQA-Eval (Critical-Thinker) 4.44 (±0.08) 4.02 (±0.13) 4.08 (±0.17) 0.711 4.64 (±0.06) 3.97 (±0.08) 4.10 (±0.08) 0.624 0.634
IQA-Eval (Adaptability-Seeker) 4.24 (±0.05) 3.67 (±0.11) 3.75 (±0.11) 0.713 4.52 (±0.08) 3.84 (±0.07) 3.84 (±0.09) 0.637 0.650
IQA-Eval (Clarity-Seeker) 4.45 (±0.07) 3.77 (±0.15) 3.80 (±0.12) 0.747 4.60 (±0.04) 3.85 (±0.04) 3.94 (±0.06) 0.676 0.690
Table 4: IQA-Eval evaluation results (helpfulness and fluency) of IQA models. Correlations are between the LEA evaluation in each row and human evaluations.
In Table 4, we show the results of model (with personas) evaluations. To accurately simulate persona distribution, each interaction is executed multiple times, with different personas (including the standard one) assigned to LEA based on their distribution proportions. This method ensures that each persona’s influence and characteristics are proportionally represented in the simulation, reflecting their respective prevalence within the overall distribution. The final score for a persona is an average of all experiment results.
Assigning “Expert” persona leads to a significant decrease in the number of queries because this “Expert” already possesses relevant knowledge and only needs key explanations. The “Clarity-Seeker” requires the highest number of interaction turns among all personas, since it requires additional clarity on questions and entities. Notably, when interacting with TextBabbage and Davinci, the “Clarity-Seeker” achieves the highest accuracy due to a clear understanding obtained from the details of the questions. The results for “Critic-Thinker” and “Adapability-Seeker” in Tables 4 and 4, which rarely improve upon the standard persona, suggest these personas are less common within the overall distributions. It matches our original expectations. In Table 4, the large variation of “# Queries” across different personas indicates that it significantly influences the interaction strategies of LEAs. Furthermore, the accuracy results are around the same level after adding the persona. It demonstrates that the overall performance doesn’t drop with the assignment of personas.
The results in the two tables show that: after assigning persona to LEAs to more accurately represent the specific group of workers, general correlation scores substantially increase. It indicates that assigning specific personas enables LEAs to operate in a more detailed and precise manner, and enhances LEAs to perform IQA-Eval in a more fine-grained and human-aligned way.
6 Benchmarking LLMs with IQA-Eval on more Types of Questions
6.1 Datasets
To evaluate the robustness and generalizability of our evaluation framework, we conduct benchmarking across different models on two distinct question answering datasets, each offering unique challenges and complexities requiring advanced reasoning. AmbigQA 35 is a collection of 14,042 annotated questions sourced from the NQ-OPEN benchmarks 22, an open-domain QA dataset. It focuses on questions with inherent ambiguities, reflecting the complexity encountered in real-world queries. These ambiguities often involve diverse aspects such as events, entity references, and answer types, resulting in multiple plausible answers for each question. HotpotQA 53 comprises 113,000 question-answer pairs sourced from Wikipedia, which require multi-hop reasoning spanning multiple documents. It contains a rich array of intricate questions that demand the synthesis of information from various texts to determine accurate answers. In this benchmark, we select 500 questions from each dataset to form a dataset containing 1,000 complex multi-hop and ambiguous questions.
6.2 LLMs to Benchmark
IQA Models HotpotQA AmbigQA
[…]
TextDavinci 4.72 4.87 1.22 0.45 - - - -
TextBabbage 4.70 4.88 1.74 0.37 - - - -
Davinci 4.27 4.52 1.68 0.32 - - - -
GPT3.5 4.72 4.95 1.49 0.63 4.91 4.97 1.89 0.60
GPT4 4.78 4.96 1.12 0.66 4.89 4.95 1.06 0.72
Claude 4.82 4.99 1.26 0.58 4.89 4.94 1.36 0.62
Llama2 4.70 4.95 1.32 0.55 4.96 4.94 1.79 0.52
Zephyr 4.64 4.88 1.01 0.40 4.38 4.66 1.03 0.45
Table 5: IQA-Eval benchmarking results on HotpotQA and AmbigQA datasets.
Apart from TextDavinci, TextBabbage and Davinci we benchmark more LLMs: GPT3.5, GPT4, Claude, Llama2 and Zephyr. The checkpoints for Llama2 and Zephyr are Llama-2-7B and Zephyr-alpha, respectively. GPT3.5 is used as LEA in our experiments.
6.3 Benchmarking Results s
The IQA evaluation benchmarks are presented in Table 5. We divide IQA Models into two categories: weak IQA Models (TextDavinci, TextBabbage, and Davinci) and strong IQA Models (GPT3.5, GPT4, Claude, Llama2, and Zephyr). Weak IQA Models can assist the LEA with answering HotpotQA questions, but due to their knowledge limitations, they cannot help much with AmbigQA questions. Zephyr achieves the lowest performance compared to other strong IQA Models. On the HotpotQA dataset, Zephyr’s accuracy performance is only comparable to the strongest one among weak IQA Models, TextDavinci. Most “Helpfulness” and “Fluency” scores are high (exceeding 3 out of 5), especially for strong IQA Models like GPT4. For the “# of queries”, it is uncommon for interactions to extend beyond two turns. As on HotpotQA, most interactions conclude at the beginning of the second turn, as IQA models have effectively guided users to reach the answers. For AmbigQA, some conversations last longer whereas LEA spends additional turns on clarifying ambiguous entities before approaching the final answer.
6.4 Analysis
IQA Models HotpotQA AmbigQA
GPT3.5 0.43 0.62
GPT4 0.46 0.63
Claude 0.28 0.5
Llama2 0.24 0.29
Zephyr 0.25 0.31
Table 6: Accuracy of IQA Models (recent LLMs) on two datasets (Non-interactive setting).
Stronger IQA models require fewer turns in interactions.
On the more challenging AmbigQA, the stronger model, GPT4, typically requires only one turn to assist LEA in solving questions with high accuracy. In contrast, less powerful models like Llama2 and GPT3.5 need more turns to clarify ambiguous entities, and their accuracy rates are lower. A similar trend is observed on the HotpotQA.
We obtain a similar model ranking with a much lower cost.
Compared to Chatbot Arena  [^2] https://huggingface.co/spaces/lmsys/chatbot-arena-leaderboard, our accuracy-based ranking of IQA Models follows a similar trend: GPT4 > Claude > GPT3.5 > Llama2 > Zephyr. In addition, our evaluation method, IQA-Eval, is fully automated. Our method makes it a cost-effective alternative for large-scale evaluations.
Evaluation of interaction performance does not always match Non-Interaction performance.
In interaction evaluations, accuracy on final results is not the only metric to show IQA Models’ performance. The quality of intermediate responses is always a significant aspect. On both “helpfulness” and “fluency” metrics, Claude is always the best IQA Model on HotpotQA questions, while on AmbigQA, Llama2 and GPT3.5 outperform GPT4. IQA Model rankings on these two aspects differ from those in Chatbot Arena (non-interaction).
The performance of IQA Models largely affects the final performance.
The accuracies in both Table 5 and Table 6 are highly correlated. The Pearson correlations of the accuracy on both datasets are 0.77 and 0.87, respectively. A strong IQA model, such as GPT4, can lead the LEA to finish tasks and largely improve the LEA’s performance on those tasks. Weak assistants may drag down the LEA’s performance, such as the performance of the LEA on both datasets decreases after interacting with Zephyr.
7 Conclusion
To conclude, we introduced IQA-Eval, a novel approach for evaluating interactive question-answering systems using large language models. Our methodology achieves automatic interaction generation and evaluation with LEA, and enhances the evaluation process by assigning personas to LEA for better matching diverse groups of people. We show that our approach aligns closely with real human interactions and judgment, indicating that a scalable, automatic IQA-Eval process can be achieved. We providing insights on recent LLM’s capability in conducting IQA with IQA-Eval  which would cost \$5,000 for human evaluations.
Appendix A Limitations
In conducting this study, certain limitations have influenced our scope and findings. Our methodology was applied exclusively to multi-choice question-answering tasks due to constraints imposed by the datasets used. Moreover, we do not investigate how allowing prompt editing could affect the results. This specialization, while focused, limits the generalizability of our findings across the wider array of question answering formats that exist in both academic research and practical applications.
We advocate for subsequent research efforts to extend the application of our proposed evaluation strategies to a more diverse set of question answering tasks, beyond the multi-choice format. Furthermore, there is a significant opportunity to test these methods with a broader spectrum of LLMs, including those at the cutting edge of the field. Such expansions would not only validate the versatility and robustness of our approaches but also potentially uncover additional insights into the nuances of LLM interaction and performance in varied contexts.
Appendix B Prompts
Both B.1 and B.2 are adapted from Lee et al. [2023]. All the following prompts are from the view of LEA in interactions.
B.1 Interaction Evaluation Prompt
We follow the setting in the data and evaluate conversations worker-wise. In other words, we combine conversations for each worker and send all of them to evaluators. The prompt we send to APIs follows the following format:
You are a helpful and precise assistant for checking the quality
of the AI assistant’s responses in interactions.
{Question 1}
{Golden Answer 1}
{Conversation 1}
{User Answer 1}
$\cdots$
{Question $n$}
{Golden Answer $n$}
{Conversation $n$}
{User Answer $n$}
Please evaluate the above conversations between user and AI assistant
by using the following metrics:
Fluency (5-point Likert): How clear (or fluent) were the responses
from the AI Assistant?
Helpfulness (5-point Likert): Independent of its fluency, how helpful
was having access to the AI Assistant compared to not having access?
Helpfulness (free-form): Why did you find the AI Assistant helpful
or unhelpful?
Please output each of the above metrics line-by-line.
B.2 Interaction Generation Prompt
Since this is a multi-choice question answering task, the full prompt for models is as follows:
You are mimicking a human.
You are trying to choose the correct answer to the given question.
Please ask an assistant sub-questions for help approaching answers.
In each turn, please only ask one sub-question to interact with an
assistant. In the sub-questions, please include all necessary
information, such as the question and options, in the original
question. If you know the answer, please output "So, the answer
is: A, B, C, or D."
{QA Question and choices}
{User Model’s query: [question 1]}
{Assistant’s answer: [answer 1]}
{User Model’s query: [question 2]}
{Assistant’s answer: [answer 2]}
$\cdots$
{User Model’s query: [question $n$]}
{Assistant’s answer: [answer $n$]}
{User Model’s final answer}
If the current turn reaches the maximum number we set, the system prompt before “{Question}” looks as follows:
Please choose the correct answer to the given question. Please
output "So, the answer is: A, B, C, or D."
B.2.1 QA Question and choices format
The question prompt for multi-choice questions in MMLU is as follows:
<question>
A. <option A>
B. <option B>
C. <option C>
D. <option D>
For HotpotQA and AmbigQA datasets, the question prompt only contains a question, such as: <question>
B.3 Persona Prompts
We design distinct prompts for each persona. Both prompts in meta-evaluation and model interaction modules change with personas.
In model interaction prompts, we only modify the first sentence based on personas. See all persona prompts in Table 7.
Persona Persona Interaction Description Persona Evaluation Description
Expert You are mimicking a knowledgeable human who can quickly understand new concepts. You need help from an assistant to learn and answer questions. The AI Assistant helps a knowledgeable human to answer a question. The assistant should provide straightforward, informative, and in-depth answers to human questions.
Critical-Thinker You are mimicking a human who prefers interactions rich in critical information. You need help from an assistant and try to get critical information from it to answer the following questions. The AI Assistant should provide clear, non-vague, and precise information or options and help user deduce answers. (Detailed evaluation criteria were indicated but not fully transcribed due to length.)
Adaptability-Seeker You are mimicking a human who prefers an adaptable assistant who can always understand his questions. You need help from an assistant to answer questions. The AI Assistant helps a human who prefers an adaptable assistant. The assistant should understand user’s questions, provide related options, and help user deduce answers.
Clarity-Seeker You are mimicking a human who prefers clear information in conversations. You need help from an assistant and want to get clear information from it to answer questions. The AI Assistant helps a human who prefers clear information in conversations. The AI should provide non-vague, precise information to help user deduce answers.
Table 7: Persona Interaction and Evaluation Descriptions
Appendix C Self-Enhancement Bias
We designed a methodology to further mitigate self-enhancement bias. In this experiment, multiple LEA models evaluate the performance of IQA models during each interaction (“Multi-perspective”). In other words, we introduced third-party evaluations, where various LEA models assess the IQA models’ performance instead of relying solely on the LEA model itself involved in the interaction. After evaluation, we use the average score from all LEA models as the final score. The results of IQA-Eval-Multi-perspective look as in Table 8. The correlations between IQA-Eval-Multi-perspective and human evaluations are in Table 9.
Helpfulness Fluency # Queries Accuracy
LEA models TDA TB DA TDA TB DA TDA TB DA TDA TB DA
Human 4.60 3.84 3.52 4.35 3.84 3.22 1.78 2.57 2.66 0.69 0.52 0.48
IQA-EVAL-GPT4-MP 4.32 3.70 3.53 4.57 3.74 3.68 1.57 2.27 2.37 0.87 0.83 0.67
IQA-EVAL-Claude-MP 3.96 3.13 3.10 4.29 3.51 3.22 2.20 2.67 2.07 0.67 0.53 0.57
IQA-EVAL-GPT3.5-MP 3.98 3.23 3.04 4.41 3.67 3.59 1.57 1.77 2.00 0.63 0.47 0.53
Table 8: IQA-EVAL-Multi-Perspective Results of IQA Models. MP indicates “Multi-Perspective”. Bold numbers indicate they are the closest to human results.
LEA models Helpfulness Fluency Overall
GPT4 0.702 0.601 0.624
Claude 0.663 0.613 0.602
GPT3.5 0.641 0.552 0.533
Table 9: Pearson Correlation ($\rho$) between IQA-EVAL-Multi-Persepctive evaluations and human judgments.
Appendix D Additional Experiments on Natural Questions
We also benchmark IQA models in another dataset called Natural Questions Kwiatkowski et al. [2019b]. This dataset comprises authentic questions posed by users about Wikipedia articles, demanding true multi-turn dialogues for resolution, akin to the setup in multi-turn conversational QA dataset QuAC Choi et al. [2018]. The experiment results are as in Tables 12 and 12. All numbers of queries in the two tables are around 3, and each response to a query from IQA models contains an average of 2 sentences.
Given the number of sentences in each IQA model’s response in Table 12, the non-interactive outputs are roughly equivalent to about two interaction turns, less than three turns in interactive outputs. Thus, The interaction process of IQA-EVAL involves not only reasoning processes but also simulating genuine interactive multi-turn conversations. This suggests that the performances shown in tables 12 and 12 above are driven more by multi-turn interactions than by reasoning processes. Furthermore, these interactions lead to enhanced accuracy, as demonstrated by the superior results in the first two tables compared to those in the last table (non-interactive).
IQA models Helpfulness Fluency # Queries Accuracy
GPT3.5 4.86 4.88 2.82 0.42
Claude 4.88 4.90 3.02 0.38
Llama2 4.90 4.84 3.18 0.34
Zephyr 4.84 4.90 3.02 0.28
Table 10: IQA-Eval benchmarking results on the Natural Questions by Claude-3
IQA models Helpfulness Fluency # Queries Accuracy
GPT3.5 4.12 5.00 2.76 0.44
Claude 4.02 5.00 2.76 0.40
Llama2 3.20 4.84 3.08 0.32
Zephyr 3.30 4.86 2.92 0.36
Table 11: IQA-Eval benchmarking results on the Natural Question by GPT-4
IQA models # Sentences Accuracy
GPT3.5 4.66 0.38
Claude 3.16 0.34
Llama2 5.21 0.30
Zephyr 4.68 0.24
Table 12: Average number of sentences and accuracy scores of IQA Models (non-interactive setting)
Appendix E Sensitive to Persona Distributions
We expect that IQA-EVAL is sensitive to incorrect persona assignments. We conduct two new experiments to study the effects of changing persona assignments. When the persona distribution is incorrect (such as 20% Expert in Table 13), the performance of IQA-EVAL shows a lower correlation with human evaluations.
Moreover, the last two lines in Table 13 describe the correlation between human evaluations and IQA-EVAL within a sub-group only containing pure experts. The correlation results in line “IQA-EVAL (Pure Expert)” represent that (1) our personas accurately represent the pure expert group, as its correlation with the line “Human (Pure Expert)” remains nearly consistent with those in line “IQA-EVAL (Expert)” and (2) given this completely correct persona distribution, our IQA-EVAL correlates well with human evaluations.
Helpfulness Fluency
LEA models TDA TB DA $\rho$ TDA TB DA $\rho$
Human 4.60 3.84 3.52 4.35 3.84 3.22
IQA-EVAL (Expert) 4.17 3.08 3.12 0.756 4.47 3.84 3.40 0.787
IQA-EVAL (20% Expert) 4.31 3.26 3.44 0.708 4.62 4.09 3.65 0.741
IQA-EVAL (40% Expert) 4.21 3.14 3.23 0.751 4.49 3.88 3.44 0.779
IQA-EVAL (60% Expert) 4.11 3.01 3.00 0.725 4.43 3.77 3.34 0.734
IQA-EVAL (80% Expert) 4.02 2.90 2.79 0.680 4.30 3.56 3.12 0.703
Human (Pure Expert) 4.69 4.00 3.73 4.36 3.96 3.26
IQA-EVAL (Pure Expert) 4.37 3.57 3.33 0.778 4.20 3.40 2.97 0.786
Table 13: IQA-EVAL results under different persona distribution on the expert persona.
Appendix F Accurate QA models are preferred by humans in IQA-EVAL
The quote from our cited paper Lee et al. [2023] is “[…] perception of helpfulness is not necessarily reflected in the overall interaction accuracy.” It describes the conclusion of multiple tasks in that paper (e.g. text summarization, social dialogue, QA). However, in the QA settings, Table 3 in Lee et al. [2023] shows that humans prefer accurate models on the QA task.
We also conducted experiments (1) using LEA to evaluate interactions between LEAs and IQA models (interactive) and (2) using LEA to evaluate direct answers generated by IQA models (non-interactive). Our experiments in Tables 15 and 15 show that LEA models prefer accurate models, which aligns well with the conclusion from human annotations.
Helpfulness Fluency Accuracy
LEA models
GPT
-3.5
Claude
-instant
Llama2
-8b
Zephyr
-Alpha
GPT
-3.5
Claude
-instant
Llama2
-8b
Zephyr
-Alpha
GPT
-3.5
Claude
-instant
Llama2
-8b
Zephyr
-Alpha
IQA-EVAL-GPT4 4.60 4.60 3.83 4.27 4.97 5.00 4.87 4.93 0.93 0.93 0.83 0.93
IQA-EVAL-Claude 4.90 5.00 4.97 4.97 4.87 5.00 4.93 4.87 0.73 0.8 0.57 0.73
Table 14: Evaluation results of interactions between LEA and IQA models.
Helpfulness Fluency Accuracy
LEA models
GPT
-3.5
Claude
-instant
Llama2
-8b
Zephyr
-Alpha
GPT
-3.5
Claude
-instant
Llama2
-8b
Zephyr
-Alpha
GPT
-3.5
Claude
-instant
Llama2
-8b
Zephyr
-Alpha
IQA-EVAL-GPT4 4.33 4.17 2.70 3.53 5.00 4.97 4.13 4.33 0.83 0.80 0.47 0.57
IQA-EVAL-Claude 4.97 5.00 4.53 4.87 4.97 4.97 4.47 4.97 0.83 0.80 0.47 0.57
Table 15: Evaluation results of non-interactions (direct answers) between LEA and IQA models.